# Multi-Lattice Sampling of Quantum Field Theories via Neural Operator-based Flows

### v3

@@ Abstract
We consider the problem of sampling discrete field configurations $\phi$ from the Boltzmann distribution $[d\phi]Z^{-1}e^{-S[\phi]}$, where $S$ is the lattice-discretization of the continuous Euclidean action $\mathcal{S}$ of some quantum field theory. Since such densities arise as the approximation of the underlying functional density $[\mathcal{D}\phi(x)]\mathcal{Z}^{-1}e^{-\mathcal{S}[\phi(x)]}$, we frame the task as an instance of operator learning. In particular, we propose to approximate a time-dependent operator $\mathcal{V}_{t}$ whose time integral provides a mapping between the functional distributions of the free theory $[\mathcal{D}\phi(x)]\mathcal{Z}_{0}^{-1}e^{-\mathcal{S}_{0}[\phi(x)]}$ and of the target theory $[\mathcal{D}\phi(x)]\mathcal{Z}^{-1}e^{-\mathcal{S}[\phi(x)]}$. Whenever a particular lattice is chosen, the operator $\mathcal{V}_{t}$ can be discretized to a finite dimensional, time-dependent vector field $V_{t}$ which in turn induces a continuous normalizing flow between finite dimensional distributions over the chosen lattice. This flow can then be trained to be a diffeormorphism between the discretized free and target theories $[d\phi]Z_{0}^{-1}e^{-S_{0}[\phi]}$, $[d\phi]Z^{-1}e^{-S[\phi]}$. We run experiments on the $\phi^{4}$-theory to explore to what extent such operator-based flow architectures generalize to lattice sizes they were not trained on and show that pretraining on smaller lattices can lead to speedup over training only a target lattice size.
@@ 1 Introduction
Consider $\mathcal{S}$ as an action characterizing a quantum field theory, with $S$ representing its discretization to a lattice. Albergo et al. [1] suggest a method for sampling from the lattice quantum field theory described by $S$. This involves a normalizing flow parameterizing a probability density $q_{\theta}(\phi)$ of discrete fields over the lattice, and optimizing the parameters $\theta$ until $q_{\theta}(\phi)$ closely approximates $\frac{e^{-S[\phi]}}{Z}$, where $Z=\int[d\phi]e^{-S[\phi]}$ is the normalizing constant.
@@
Operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and model should capture the underlying continuous physics. Kovachki et al. [2] term this property of models discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather that of convergence. In this work, we apply the same idea to the task of sampling from lattice quantum field theories, motivated by the fact that lattice field theories also emerge as the discretization of continuous field theories.
@@
Suppose now that the field theory is defined on some domain $D$. Once a lattice, as a discretization of $D$, is chosen, one can construct a continuous normalizing flow [3] as a time-dependent vector field $V_{t}$ that parametrizes the direction along which probability mass moves. Generalizing this idea, we propose to parametrize a time-dependent operator $\mathcal{V}_{t}$ from the space of functions on $D$ to itself that defines the direction in which functional probability mass moves. Such an operator can then be used to map the functional distributions $[\mathcal{D}\phi(x)]\mathcal{Z}_{0}^{-1}e^{-S_{0}[\phi(x)]},[\mathcal{D}\phi(x)]\mathcal{Z}_{1}^{-1}e^{-S_{1}[\phi(x)]}$ to one another. Computationally the operator $\mathcal{V}_{t}$ can only be accessed by a choice of a lattice which induces a vector field $V_{t}$ as the discretization of $\mathcal{V}_{t}$. We then train this vector field to be a diffeomorphism between the discretized free and target theories, $[d\phi]Z_{0}^{-1}e^{-S_{0}[\phi]}$ and $[d\phi]Z_{1}^{-1}e^{-S_{1}[\phi]}$. The upside of using a operator-based flow will be that a single model can be used to operate on multiple discretizations of the same underlying continuous system. Figure 1 provides a schematic overview of the objects and their relation in this paragraph.
{tikzcd}
@@
Figure 1: Schematic overview of the probability distributions of interest. The top row shows the functional distributions of the free theory and the target theory connected by the time dependent operator $\mathcal{V}_{t}$. Descending to the bottom row means approximating all the objects of the top row on a lattice. In particular, in the bottom row all objects are finite dimensional, well-defined and can be numerically worked with.
The structure of the paper is as follows: Section §2 contains the relevant background on continuous normalizing flows, Boltzmann densities, the $\phi^{4}$ quantum field theory and neural operators. In Section §3 we describe an operator-based normalizing flow architecture that we used in our experiments. Section §4 documents the experiments on the $\phi^{4}$ theory.
@@ 2.1 Continuous Normalizing Flows
A continuous normalizing flow [3] is a density estimator that operates by pushing forward a simple, usually Gaussian, initial density $q_{0}$ along a parametric, time-dependent vector field $V_{\theta}:[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Explicitly, the pushforward density $q_{\theta}$ is given by
@@
where $\nabla$ is the divergence operator in the spatial coordinates and $x_{t}$ is the integral curve of $V_{\theta}$ that passes through $x_{1}$ at $t=1$. In this work all normalizing flows will be continuous normalizing flows, and we will refer to them as normalizing flows or even just flows for brevity.
@@
The Boltzmann distribution of an energy function [^2] Assuming that $\exp(-f)$ is integrable. $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a probability distribution with density function
$p(x)=\frac{1}{Z}e^{-f(x)}$ (2)
@@
$KL[q_{\theta},p]$ $=\mathbb{E}_{x\sim q_{\theta}}\left[\log q_{\theta}(x)-\log p(x)\right]$ (3)
$=\mathbb{E}_{x\sim q_{\theta}}\left[\log q_{\theta}(x)+f(x)\right]+Z$ (4)
where $q_{\theta}$ is the density realized by the normalizing flow (1). Once a density $q_{\theta}$, approximating $p=Z^{-1}e^{-f}$, is learnt, one can use importance sampling to correct for small inaccuracies of $q_{\theta}$ when estimating the expected value of an observable $\mathcal{O}$
$\langle\mathcal{O}\rangle:=\mathbb{E}_{x\sim p}\left[\mathcal{O}(x)\right]=\mathbb{E}_{x\sim q_{\theta}}\left[\mathcal{O}(x)\frac{p(x)}{q_{\theta}(x)}\right]$ (5)
2.2 The $\phi^{4}$ (Lattice) Quantum Field Theory [^3] We recommend the book [17, Chapter 15] for further details on lattice field theories.
Let us now consider the Euclidean action on real valued scalar fields $\phi(x)$ with periodic boundary conditions on the $D$-dimensional hypercube of edge length $L$, $\phi:(\mathbb{R}/L\mathbb{Z})^{D}\rightarrow\mathbb{R}$, for some constants $m^{2}$ and $g$
$\mathcal{S}[\phi]$ $=\int_{(\mathbb{R}/L\mathbb{Z})^{D}}d^{D}x\left[(\nabla\phi)^{2}+m^{2}\phi^{2}+g\phi^{4}\right]$ (6)
where we dropped the argument $x$ of the field $\phi(x)$ for notational convenience. To estimate the expectation value of an observable $\mathcal{O}$, we need to average over all field configurations that satisfy the boundary conditions, with each configuration weighted by the exponential of the negative action
@@
The action $\mathcal{S}$ corresponds to the energy function $f$ of a Boltzmann density and the denominator $\mathcal{Z}=\int\mathcal{D}\phi\,e^{-\mathcal{S}[\phi]}$ to the normalizing constant as introduced in Section §2.1.
Equations (6) and (7) describe an infinite dimensional system. To tackle it numerically, one first needs to discretize it to a lattice. This comes at the cost of losing the information contained in the high-frequency components as the highest possible frequency of a periodic function on a lattice with edge length $L$ with $N$ nodes is $\frac{2\pi N}{L}$. The hope is that one can do the same on larger and larger lattices, and as the lattice approaches the continuum limit, the error due to discretization converges to zero.
@@
After these substitutions we end up with the following discretised action on the lattice,
@@ 2.3 Neural Operators
Neural operators[2] are trainable function-to-function mappings. In particular, both their domains and codomains are infinite dimensional function spaces. In practice, one works with neural operators by choosing a mesh/lattice $X\subset\mathbb{R}^{n}$, representing functions by their evaluations on $X$ and let the neural operator operate on this collection of evaluations. By design, neural operators can be evaluated on lattices of different size. Importantly, if a neural operator is applied to a sequence of meshes $X_{i}$, approaching the continuum limit $X_{i}\rightarrow\mathbb{R}^{n}$, it converges to the underlying continuous operator. The main use case of neural operators is to approximate the solution of partial differential equations, i.e. learn the mapping from an initial condition to the time evolved state after some time $\Delta t$ (Figure 2), but have also been applied for multi-resolution generative modelling [18, 19]. We will use them for parametrizing a flow, i.e. a vector field $\mathcal{V}_{t}$ connecting the free theory (base density) to the $\phi^{4}$-theory (target density) in a way that can be evaluated at any mesh.
Figure 2: An operator that maps an initial condition $u(0,x)$ (top row) to its time-evolved state $u(\Delta t,x)$ (bottom row). The time evolution is given by the heat equation $\Delta u=\partial_{t}u$. The blue dots denote the evaluation of $u(0,x)$ on a discrete mesh, while the orange dots denote the output of the operator (a convolution in this case) evaluated on that same mesh. As the mesh gets denser, the operator becomes a better approximation of the map between the continuous $u(0,x)$ (blue curve) and $u(\Delta t,x)$ (orange curve).

### v4

@@ Abstract
We consider the problem of sampling lattice field configurations on a lattice from the Boltzmann distribution corresponding to some action. Since such densities arise as approximationw of an underlying functional density, we frame the task as an instance of operator learning. We propose to approximate a time-dependent neural operator whose time integral provides a mapping between the functional distributions of the free and target theories. Once a particular lattice is chosen, the neural operator can be discretized to a finite-dimensional, time-dependent vector field which in turn induces a continuous normalizing flow between finite dimensional distributions over the chosen lattice. This flow can then be trained to be a diffeormorphism between the discretized free and target theories on the chosen lattice, and, by construction, can be evaluated on different discretizations of spacetime. We experimentally validate the proposal on the 2-dimensional $\phi^{4}$-theory to explore to what extent such operator-based flow architectures generalize to lattice sizes they were not trained on, and show that pretraining on smaller lattices can lead to a speedup over training directly on the target lattice size.
@@ 1 Introduction
Consider an action $\mathcal{S}$, characterizing a quantum field theory, and $S$, a discretized representation of $\mathcal{S}$ on a lattice. Albergo et al. [1] suggest a method for sampling from the lattice quantum field theory described by $S$ by using a normalizing flow parameterizing a probability density $q_{\theta}(\phi)$ of discrete fields over the lattice, and optimizing the parameters $\theta$ until $q_{\theta}(\phi)$ closely approximates $\frac{e^{-S[\phi]}}{Z}$, where $Z=\int[d\phi]e^{-S[\phi]}$ is the normalizing constant.
@@
Operator learning promotes the viewpoint that the lattice/mesh is merely a computational tool, and the model should capture the underlying continuous physics. Kovachki et al. [2] refer to this property as discretization invariance.  [^1] Discretization invariance means that the neural operator evaluated on finer and finer discretizations approximates the continuous operator. Thus, strictly speaking, it is not a requirement of invariance rather one of convergence. In this work, we apply the same idea to the task of sampling from lattice quantum field theories, motivated by the fact that lattice field theories also emerge as the discretization of continuous field theories.
(Lattice) Quantum Field Theory
A quantum field theory is defined by an action functional $\mathcal{S}$ that assigns a scalar to field configurations on some domain, typically denoted by $\phi(x)$, where $x$ represents spacetime coordinates. After a Wick rotation (a transformation that makes time act like a spatial dimension), the problem of interest is to sample from the functional density $\frac{1}{\mathcal{Z}}e^{-\mathcal{S}[\phi]}$, where $\mathcal{Z}$ is a normalizing constant, often called the partition function. This density represents the probability distribution over all possible field configurations in Euclidean spacetime. The goal is to understand the behavior of the system by computing expected values of observables (like correlation functions) with respect to $\frac{1}{\mathcal{Z}}e^{-\mathcal{S}[\phi]}$.
In quantum field theory, an observable $\mathcal{O}$, is simply a function of the field configuration $\phi(x)$. For example, observables could include the average field value, or correlations between field values at different points in space. When working with the functional density $\frac{1}{\mathcal{Z}}e^{-\mathcal{S}[\phi]}$, the objective is to compute the expected values of these observables with respect to this distribution, i.e. quantities of the form
$\langle O\rangle=\int\mathcal{D}\phi\,O[\phi]\,\frac{1}{\mathcal{Z}}e^{-\mathcal{S}[\phi]},$ (1)
This integral represents the average of $O$ across all possible field configurations, weighted by the probability density $\frac{1}{\mathcal{Z}}e^{-\mathcal{S}[\phi]}$. These expected values correspond to measurable, physical quantities in experiments, as they represent the statistical average over many possible configurations of the quantum field. In practice, we cannot measure individual field configurations; instead, we observe properties like particle masses, energy densities, and correlation functions, each of which can be computed as the expected value of an appropriate observable. Therefore, expected values provide the essential link between theoretical calculations and physical measurements, making them central to understanding the field’s behavior.
To do this computationally, quantum field theories are typically discretized by approximating the continuous spacetime domain with a lattice structure. This discretization translates the continuous field $\phi(x)$ into a set of field values $\phi_{i}$ defined at specific lattice points. By introducing a finite lattice spacing $a$, we can approximate the continuous integrals over spacetime by discrete sums over lattice points, making the model tractable for numerical simulations. However, discretization inherently introduces approximations. For instance, the lattice spacing $a$ imposes a high-frequency cutoff, as fluctuations with wavelengths smaller than $a$ (frequencies above $1/a$) cannot be captured on the lattice. Additionally, the lattice domain is typically chosen to be periodic to reduce edge effects, effectively wrapping around in each dimension. The size of this periodic domain must be sufficiently large to avoid finite-size effects, where the limited domain size influences physical quantities, potentially distorting results.
While discretization makes the theory computationally feasible, the ultimate goal is to recover the properties of the original continuous quantum field theory. To achieve this, one must take the continuum limit, in which the lattice spacing $a$ is taken to zero. In this limit, the lattice approximation is refined such that finer and finer details of the field configurations are captured, and high-frequency fluctuations, initially filtered out by the discrete lattice, are restored. As $a\to 0$, the discretized sums over lattice points approach the continuous integrals that define the original theory, and the lattice version of the expected values $\langle O\rangle$ should converge to their continuous counterparts. In practice, this requires computing the expected values of observables at several different lattice spacings and then extrapolating to estimate their values as $a\to 0$. This extrapolation process allows us to obtain results that more accurately reflect the true, measurable quantities in the continuous theory. The continuum limit is essential because only in this limit do the results accurately reflect the underlying physics without artifacts from discretization. However, taking this limit is computationally demanding, as it requires not only reducing $a$ but also increasing the number of lattice points to maintain the physical size of the domain, thus ensuring that finite-size effects remain negligible.
In this work, we address this last aspect, the issue of taking the continuum limit. Traditionally, this process requires computing observables at various lattice spacings and then extrapolating to $a\to 0$. We propose an alternative approach by training a single model that can be used across different resolutions. This model leverages the concept of neural operators, which are designed to learn continuous operators between function spaces. This approach has the potential to simplify the extrapolation process, as the trained model can provide expected values for any lattice spacing, allowing for a more continuous mapping of observables as the continuum limit is approached. Consequently, we can obtain a more accurate representation of the physical quantities in the continuum theory while mitigating the computational burden typically associated with conventional, single-lattice methods.
It is important to note that this approach is related to the renormalization group (RG) methods, which also seek to understand how physical quantities change with scale. However, rather than renormalizing the action, we focus on discretizing the action at different resolutions and aim to learn a continuous representation of the underlying field theory. Our goal is to create a model that captures the essential features of the quantum field across different scales, thus enabling a more direct pathway to the continuum limit.
@@
Suppose now that the field theory is defined on some domain $D$, i.e. fields are functions living on $D$. We can reduce this system to a finite dimensional one by introducing a lattice, i.e. a collection of regularly spaced points, in $D$, and model fields as functions on this finite collection of lattice sites. In this setting, a continuous normalizing flow [3] can be introduced as a time-dependent vector field $V_{t}$ that parametrizes the direction along which probability mass moves. Generalizing this idea, we propose to parametrize a time-dependent operator $\mathcal{V}_{t}$ from the space of functions on $D$ to itself that defines the direction in which functional probability mass moves. Such an operator can then be used to map the functional distributions $[\mathcal{D}\phi(x)]\mathcal{Z}_{0}^{-1}e^{-\mathcal{S}_{0}[\phi(x)]},[\mathcal{D}\phi(x)]\mathcal{Z}_{1}^{-1}e^{-\mathcal{S}_{1}[\phi(x)]}$, with corresponding actions $\mathcal{S}_{0},\mathcal{S}_{1}$, to one another. Computationally the operator $\mathcal{V}_{t}$ can only be accessed by a choice of a lattice which induces a vector field $V_{t}$ as the discretization of $\mathcal{V}_{t}$. We then train this vector field to be a diffeomorphism between the discretized free and target theories, $[d\phi]Z_{0}^{-1}e^{-S_{0}[\phi]}$ and $[d\phi]Z_{1}^{-1}e^{-S_{1}[\phi]}$. The upside of using an operator-based flow will be that a single model can be used to operate on multiple discretizations of the same underlying continuous system. Figure 1 provides a schematic overview of the objects and their relation in this paragraph.
[…]
@@
Figure 1: Schematic overview of the probability distributions of interest. The top row shows the functional distributions of the free theory and the target theory connected by the time dependent operator $\mathcal{V}_{t}$. Moving to the bottom row corresponds to approximating each object from the top row on a discrete lattice. In particular, in the bottom row all objects are finite dimensional, well-defined and can be worked with numerically.
Related Work
Albergo et al. [1] proposed the use of normalizing flows for sampling from lattice field theories. Since then, a variety of methods have emerged for sampling scalar [4, 5, 6, 7, 8, 9, 10, 11, 12] as well as gauge field theories [13, 14, 15, 16, 17, 18, 19]. As the action of the theory is available, one can train these normalizing flow models in a data-free setting, using only evaluations of the action and its gradients as the training signal. The simplest way of doing this is to rely on the reverse Kullback–Leibler (KL) divergence, $KL(q_{\theta},p)$, between the density represented by the flow $q_{\theta}$ and the target Boltzmann density $p(\phi)=e^{-S(\phi)}/Z$. This approach encourages the learned density $q_{\theta}$ to match the high-probability regions of the target distribution, which can be an effective strategy for capturing prominent features of $p$. However, minimizing the reverse KL divergence can lead to mode collapse [12, 20, 21, 22, 23], meaning that the model density $q_{\theta}$ fails to represent the full diversity of the target distribution by focusing too narrowly on dominant modes. This limitation has motivated recent work to explore alternative training objectives that mitigate mode collapse. Vaitl et al. [20, 21, 22] introduce improved estimators of the KL divergence. A different objective motivated by the continuity equation describing the flow of probability mass between the target and prior was considered in [23].
Most of these approaches focus primarily on training models for a single lattice size, which limits their adaptability across different lattice spacings. Notably, Gerdes et al. [7] devise a method to embed a discrete convolutional kernel into a kernel for a larger lattice. In contrast, our approach treats the kernels as inherently continuous objects from the outset, enabling us to work with their discretizations across multiple lattice sizes.
@@ 2.1 Continuous Normalizing Flows
A continuous normalizing flow [3] is a density estimator and sampler that operates by pushing forward a simple, usually Gaussian, initial density $q_{0}$ along a parametric, time-dependent vector field $V_{\theta}:[0,1]\times\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Explicitly, the pushforward density $q_{\theta}$ is given by
@@
where $\nabla$ is the divergence operator in the spatial coordinates and $x_{t}$ is the integral curve of $V_{\theta}$ that passes through $x_{1}$ at $t=1$. In this work, all normalizing flows will be continuous, and we will refer to them as normalizing flows or even just flows for brevity.
@@
The Boltzmann distribution of an energy function [^2] Assuming that $\exp(-S)$ is integrable. $S:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a probability distribution with density
$p(x)=\frac{1}{Z}e^{-S(x)}$ (3)
@@
$KL(q_{\theta},p)$ $=\mathbb{E}_{x\sim q_{\theta}}\left[\log q_{\theta}(x)-\log p(x)\right]$ (4)
$=\mathbb{E}_{x\sim q_{\theta}}\left[\log q_{\theta}(x)+f(x)\right]+\log Z$ (5)
where $q_{\theta}$ is the density realized by the normalizing flow (2). Once a density $q_{\theta}$, approximating $p=Z^{-1}e^{-f}$, is learnt, one can use importance sampling to correct for small inaccuracies of $q_{\theta}$ when estimating the expected value of an observable $\mathcal{O}$, i.e. a real-valued function of the configurations,
$\langle\mathcal{O}\rangle:=\mathbb{E}_{\phi\sim p}\left[\mathcal{O}(\phi)\right]=\mathbb{E}_{\phi\sim q_{\theta}}\left[\mathcal{O}(\phi)\frac{p(\phi)}{q_{\theta}(\phi)}\right]$ (6)
2.2 The $\phi^{4}$ (Lattice) Quantum Field Theory [^3] We recommend the book [24, Chapter 15] for further details on lattice field theories.
Let us now consider the Euclidean action on real-valued, continuous scalar fields $\phi(x)$ with periodic boundary conditions on the $D$-dimensional hypercube of edge length $L$, $\phi:(\mathbb{R}/L\mathbb{Z})^{D}\rightarrow\mathbb{R}$, for some constants $m^{2}$ and $g$
$\mathcal{S}[\phi]$ $=\int_{(\mathbb{R}/L\mathbb{Z})^{D}}d^{D}x\left[(\nabla\phi(x))^{2}+m^{2}\phi(x)^{2}+g\phi(x)^{4}\right]$ (7)
For the rest of this work, we will drop the the argument of $\phi(x)$ and denote fields by $\phi$ to unclutter notation. To estimate the expectation value of an observable $\mathcal{O}$, we need to average over all field configurations that satisfy the boundary conditions, with each configuration weighted by its Boltzmann weight
@@
Equations (7) and (8) describe an infinite dimensional system, where microstates are functions on $(\mathbb{R}/L\mathbb{Z})^{D}$. To tackle it numerically, one first needs to discretize the domain $(\mathbb{R}/L\mathbb{Z})^{D}$, the action $\mathcal{S}$, and fields $\phi(x)$ to a lattice. This comes at the cost of losing the information contained in the high-frequency components as the highest possible frequency of a periodic function on a lattice with edge length $L$ with $N$ nodes is $\frac{2\pi N}{L}$. The hope is that one can do the same on larger and larger lattices, and as the lattice approaches the continuum limit, the error due to discretization converges to zero.
@@
where $\mu_{i}$ denotes the generator of the lattice along the $i-$th coordinate axis, i.e. the displacement vector along neighboring lattice sites. After these substitutions we end up with the following discretised action on the lattice,
@@ 2.3 Neural Operators
Neural Operators [2] are trainable function-to-function mappings, both their domains and codomains are infinite dimensional function spaces. In practice, one works with neural operators by choosing a mesh/lattice $X\subset\mathbb{R}^{n}$, representing functions by their evaluations on $X$ and let the neural operator operate on this collection of evaluations. By design, neural operators can be evaluated on lattices of different size. Importantly, if a neural operator is applied to a sequence of meshes $X_{i}$, approaching the continuum limit $X_{i}\rightarrow\mathbb{R}^{n}$, it converges to the underlying continuous operator. The main use case of neural operators is to approximate the solution of partial differential equations, i.e. learn the mapping from an initial condition to the time evolved state after some time $\Delta t$ (Figure 2), but have also been applied for multi-resolution generative modelling [25, 26]. We will use them for parametrizing a flow, i.e. a vector field $\mathcal{V}_{t}$ connecting the free theory (base density) to the $\phi^{4}$-theory (target density) in a way that can be evaluated at any mesh.
Figure 2: An operator that maps an initial condition $u(0,x)$ (top row) to its time-evolved state $u(\Delta t,x)$ (bottom row), where the time evolution is given by the heat equation $\Delta u=\partial_{t}u$. The blue dots denote the evaluation of $u(0,x)$ on a discrete mesh, while the orange dots denote the output of the operator (a convolution in this case) evaluated on that same mesh. As the mesh gets denser, the operator becomes a better approximation of the map between the continuous $u(0,x)$ (blue curve) and $u(\Delta t,x)$ (orange curve). For our application, we will be interested in the time evolution of probability density of fields (corresponding to $u$ in the plots), along the time interval $[0,1]$ connecting the free theory to a interacting theory.